NEAR INFRARED SPECTROSCOPY OF DEUTERO-BROMOMETHYLENE

\underline{ANDREW J. MARR} and TREVOR J. SEARS, Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973-5000.

The Ã^{1}A′′ (0,1,0) ← X̃^{1}A′ (0,0,0) electronic band system of DCBr has been measured in the region 11 720 cm^{−1} to 11 780 cm^{−1}. Spectra were recorded using a transient absorption frequency modulation Ti:Sapphire spectrometer^{a} incorporating a multipass Herriott-type cell for increased sensitivity.^{b} The (K = 0 ← K = 1) and (K = 1 ← K = 0) sub-bands of both isotopomers, DC^{81}Br and DC^{79}Br, have been assigned and spectroscopic parameters determined. Transitions involving levels from higher K states are visible but their assignments are complicated by a variety of perturbations (vibronic, spin-orbit, coriolis, fermi resonance) involving the two singlet states and a background triplet state. Acknowledgements: The work was carried out at Brookhaven National Laboratory under Contract No. DE-AC02- 76CH00016 with the U.S. Department of Energy and supported by its Division of Chemical Sciences, Office of Basic Energy Sciences.

aJ. C. Bloch, R. W. Field, G. E. Hall, and T. J. Sears, J. Chem. Phys. \underline{101}, 1717 (1994)

bC. Fockenberg, A. J. Marr, T. J. Sears, and B-C. Chang, J. Mol. Spectrosc. in press (1998)